A 1.052g sample of an unknown compound, composed only of carbon and hydrogen, produced 3.19g of CO2 and 1.63g of H2O in a combustion analysis. What is the mass percent composition of hydrogen of the unknown compound?
1
Expert's answer
2019-10-01T06:03:26-0400

1.63/18 x 2= 0.181111111 g of H

3.19/44 x 12= 0.87 g of C.

0.1811111+0.87= 1.052 g of unknown compound.

0.181111/1.052 x 100%= 17.3003802% of H.